# Sensitive parameter dependence of the self-similar dynamics in the bubble breakup in a confined geometry

### v5

@@ II.6 Identical shapes translating in time
Figure 5: Translated shape functions for $e=1.0,1.5,2.0$, and 0.5 mm on linear scales in the period of good collapse [(a1) to (d1)] and on log-log scales for $-21$ to $-30$ ms with a fitting line obtained by fitting the data in the region between the two black crosses for $-21$ ms [(a2) to (d2)]. For all cases, $(R,D_{0},\nu)=(10,1,100)$ in mm or cS.
@@ II.7 Closed equation describing the shape
is independent of spacial coordinates. From this, we obtain the spacially constant velocity as the speed of the constriction point:
@@ III Discussion
III.1 Universality and separation of scales
We show in Fig. 4 that the shape dynamics in the present parameter range is self-similar in limited ranges of time and space by demonstrating the collapse onto a master curve $\Gamma(\xi)$ by rescaling specified in Eq. (6), where $\Gamma(\xi)$ is dependent on $e$. From the viewpoint of separation of scales, we expect the present $\Gamma(\xi)$ is independent from $R$, as supported by the data labeled by $R^{\ast}$ in Fig. 4 (a3). This is because among the independent characteristic lengths $e$, $D_{0}$, and $R$, $R$ ($\sim 1$ cm) is the largest scale and thus is well separated from the characteristic scales for the self-similar dynamics. We may also expect $\Gamma(\xi)$ is independent from $\eta$ because the master curve $\Gamma(\xi)$ in the previous studies [10, 11] is always independent from $\eta$. In fact, we confirmed the master curve obtained for $(e,R,D_{0},\nu)=(0.5,10,1.5,1000)$ well collapses on the master curve in Fig. 4 (b3) obtained for $(e,R,D_{0},\nu)=(0.5,10,1.0,100)$ near the constriction point (in the region around $z/z_{m}=0$ to 1.4), which suggests the independence of $\Gamma(\xi)$ from $\eta$ and $D_{0}$. The independence from $D_{0}$ of $\Gamma(\xi)$ suggested here is also natural from the viewpoint of separation of scales since $D_{0}$ is the second largest scale. The independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ suggested here is in accord with the dimensional analysis given below.
III.2 Comparison with the bubble breakup in non-confined geometry
The translation of the interface in time with keeping the same shape observed in the present study is a property shared by the breakup of a bubble in non-confined geometry as explained below. Under no confinement, the breakup occurs with the axisymmetry. Accordingly, the shape can be described in the cylindrical coordinate $(r,z,\theta)$ by $r=h(t,z)$, where the shape function $h(t,z)$ is independent of the $\theta$ coordinate from the axisymmetry. When a high viscosity $\eta$ of the surrounding fluid competes with capillary characterized by the surface tension $\gamma$, the dynamics is known to be described by the following equation (e.g., see Sec. 9.3.1 of [2]):
@@ III.3 Dimensional analysis
Here, we may expect that near the breakup point where $h_{m}\simeq z_{m}$ is small so that $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ eventually approach towards infinity. This implies that if the function $\Xi$ remains finite and non-zero, it becomes independent of $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}$ to recover Eq. (6) with Eqs. (1) and (4) at $e=0.5$ mm.
If we assume Eq. (18) is valid even for $e=1.0$ to 2.0 mm, we understood the remaining dependence of the master curve $\Gamma$ on a small length scale $e$ (and possibly on $D_{0}$, which is comparable to $e$) from separation of scales. In the time domain in which the collapse is observed in Fig. 4, while $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ are large enough, $e/z_{m}$ is not large enough so that the $e$-dependence remains in $\Gamma$. The present dimensional argument thus suggests the independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ as announced above. However, to justify Eq. (18) even for $e=1.0$ to 2.0 mm, we need to know the parameter dependence of the coefficients $c_{1}$ and $c_{2}$ in Eq. (2), which we will explore in a future study.
III.4 Analogy with critical phenomena
@@
The power law is observed for different quantities. At $H=0$, the specific heat $C$ and the susceptibility $\chi$ behave as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$, while the magnetization $M$ behaves as $M\simeq H^{1/\delta}$ at $T=T_{c}$. The exponents such as $\alpha$ and $\beta$ are called the critical exponents and their values are governed by dimensionality $d$ and symmetry characterized by the number of components $n$ of the vector representing the order parameter. In other words, the exponents do not depend on some details such as the strength of the microscopic interaction and the structure of lattice. As a result, various materials characterized by the same $n$ and $d$ share the same critical exponents, forming a universality class. Namely, the same set of critical exponents universally describes critical behaviors such as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$ of different materials, if they belong to the same universality class although some quantities such as the critical temperature $T_{c}$ are material-dependent. Furthermore, the critical exponents satisfy scaling relations such as $\alpha+2\beta+\gamma=2$ and such relations are universally satisfied by the critical exponents of any universality classes. Such remarkable features can be understood based on the scaling hypothesis as explained in Appendix, and the scaling hypothesis naturally emerges in the renormalization group theory [12, 13].
@@
which bears a strong resemblance with the self-similar structure in Eq. (6) with Eq. (2). Time $t$ ($=t-t_{c}$) and position $z$ ($=z-z_{c}$) correspond to temperature $T$ and external field $H$, respectively, for ”the order parameter $h(t,z)$,” a quantity which is non-zero before $t_{c}$ but is zero after $t_{c}$ at $z=0$: the pre-breakup dynamics of our focus corresponds to the temperature range below $T_{c}$. The master curve $\Gamma(\xi)$ corresponds to the function $\Psi(\xi)$, which is called the scaling function and defines the critical exponent $\delta$ through $\Psi(\xi)\simeq\xi^{1/\delta}$ for $\xi\gg 1$ (and thus $\delta$ in Eq. (7) corresponds to the critical exponent $\delta$). The scaling exponents $\beta$ and $\Delta$ in Eq. (2) correspond to $\beta$ and $\Delta$ in Eq. (19), where the exponents $\beta$, $\Delta$, and $\delta$ in critical phenomena satisfy Eq. (10) as shown in Appendix.
In other words, the critical exponents $\beta$, $\Delta$, and $\delta$ in the present case are the ones summarized in Tab. 1. This table suggests that these three exponents are independent, where Eq. (10) is more violated as $e$ increases, and that that the exponents are monotonic functions of the thickness $e$. In addition, we may expect from the viewpoint of separation of scales and the dimensional analysis discussed above the exponents are universally shared by the systems with different $R$ and $\eta$ for a given $e$ (at $D_{0}=1$ mm). Namely, the present universality class may be determined by a continuous parameter $e$. This is in contrast with typical critical phenomena, in which universality class is determined by discrete parameters such as $n$ and $d$, although examples of critical exponents depending on continuous parameters are known when marginal operators play a role, which includes non-linear diffusion equations [15]. In the present case, a renormalization group analysis suggests a different scenario, which will be discussed elsewhere.
@@ IV Conclusion
We investigate the pre-breakup dynamics of a bubble surrounded by a viscous liquid in a confined geometry focusing on the dependence on the lubricating film thickness $e$ and find that the dynamics is self-similar and that the scaling exponents characterizing the self-similarity is sensitively dependent on $e$. We elucidate similarity between the present singular dynamics and critical phenomena, showing that the counterpart of the critical exponents are dependent on a continuous variable $e$ and could be shared by a universality class formed by the systems with different $R$ and $\eta$ for a fixed $e$.
The analogy between the singular dynamics and critical phenomena elucidated in the present study underscores the importance of exploring symmetry in the future study of the singular dynamics by exploiting confined geometries for the following reasons. (1) In critical phenomena, which has propelled the development of the modern physics from soft and hard condensed matter to non-equilibrium systems and active matter [16, 17, 18, 19], dimensionality and symmetry have played a vital role in its development. (2) The singular dynamics is widely observed in nature and a variety of confined geometries are increasingly important in industrial applications and natural phenomena, from microfluidics [20, 21], petroleum industry [22], to geology [23]. In fact, the analogy with critical phenomena can be deepened further by developing a renormalization group analysis for the present singular dynamics, which will be discussed elsewhere.
@@ Acknowledgements.
Appendix A Scaling relations in critical phenomena
Figure 6: Snapshots obtained at 2000 fps near the detachment point for $e=0.5$ mm (top) and $e=1.0$ mm (bottom) at $t-t_{d}=-1.5$ to $+1.5$ ms, respectively. The snapshots at the center correspond to $t-t_{d}=-0$ ms, where ”the order parameter” $h_{m}$ is not precisely zero.

### v6

@@ II.7 Closed equation describing the shape
is independent of spacial coordinates. This expression means that the spacially constant velocity should be equal to the speed of the constriction point:
@@ III Discussion
III.1 Comparison with the bubble breakup in non-confined geometry
The translation of the interface in time with keeping the same shape observed in the present study is a property shared by the breakup of a bubble in non-confined geometry as explained below. Under no confinement, the breakup occurs with the axisymmetry. Accordingly, the shape can be described in the cylindrical coordinate $(r,z,\theta)$ by $r=h(t,z)$, where the shape function $h(t,z)$ is independent of the $\theta$ coordinate from the axisymmetry. When a high viscosity $\eta$ of the surrounding fluid competes with capillarity characterized by the surface tension $\gamma$, the dynamics is known to be described by the following equation (e.g., see Sec. 9.3.1 of [4]):
@@ III.2 Dimensional analysis
Here, we may expect that near the breakup point where $h_{m}\simeq z_{m}$ is small so that $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ eventually approach towards infinity. This implies that if the function $\Xi$ remains finite and non-zero, it becomes independent of $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}$ to recover Eq. (6) with Eqs. (1) and (4) at $e=0.5$ mm. In this way, we can naturally understand the self-similar structure at $e=0.5$ mm based on a dimensional argument.
In the case of $e=1.0$ to 2.0 mm, in order to justify a similar argument, different from the case of $e=0.5$ mm, we have yet to know the parameter dependence of the coefficients $c_{1}$ and $c_{2}$ in Eq. (2). Such a study would be interesting especially in terms of universality and separation of scales. For example, if Eq. (18) is still valid even for $e=1.0$ to 2.0 mm, we may understand universality of $\Gamma$ for change in $R$ but not in $e$ indicated in the present experiment: In the time domain in which the collapse is observed in Fig. 4, while $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ are large enough compared with the length scale over which the collapse is observed, $e/z_{m}$ is not large enough so that the $e$-dependence remains in $\Gamma$. In other words, the dependence on length scales could disappear if the scale is well separated from the length scale of collapse, but could remain if not.
This issue of universality and separation of scales is explored for the post-breakup dynamics [18]. Similar study for the present pre-detachment dynamics would be an interesting future problem, although this requires careful experiments. Note that experiments on the pre-detachment is technically more challenging because the pre-detachment dynamics tends to be more sensitively affected by the entry conditions.
We remark that the above tentative argument for $e=1.0$ to 2.0 mm suggests universality of $\Gamma$ for change in $\eta$. In fact, the master curve in the previous studies on the pre-detachment dynamics [16, 17] is always independent of $\eta$.
III.3 Analogy with critical phenomena
@@
The power law is observed for different quantities. At $H=0$, the specific heat $C$ and the susceptibility $\chi$ behave as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$, while the magnetization $M$ behaves as $M\simeq H^{1/\delta}$ at $T=T_{c}$. The exponents such as $\alpha$ and $\beta$ are called the critical exponents and their values are governed by dimensionality $d$ and symmetry characterized by the number of components $n$ of the vector representing the order parameter. In other words, the exponents do not depend on some details such as the strength of the microscopic interaction and the structure of lattice. As a result, various materials characterized by the same $n$ and $d$ share the same critical exponents, forming a universality class. Namely, the same set of critical exponents universally describes critical behaviors such as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$ of different materials, if they belong to the same universality class although some quantities such as the critical temperature $T_{c}$ are material-dependent. Furthermore, the critical exponents satisfy scaling relations such as $\alpha+2\beta+\gamma=2$ and such relations are universally satisfied by the critical exponents of any universality classes. Such remarkable features can be understood based on the scaling hypothesis as explained in Appendix B, and the scaling hypothesis naturally emerges in the renormalization group theory [19, 20].
@@
which bears a strong resemblance with the self-similar structure in Eq. (6) with Eq. (2). Time $t$ ($t-t_{c}$) and position $z$ ($z-z_{c}$) correspond to temperature $T$ and external field $H$, respectively, for ”the order parameter h(t,z)h(t,z),” a quantity which is non-zero before $t_{c}$ but is zero after $t_{c}$ at $z=0$: the pre-detachment dynamics of our focus corresponds to the temperature range below $T_{c}$. The master curve $\Gamma(\xi)$ corresponds to the function $\Psi(\xi)$, which is called the scaling function and defines the critical exponent $\delta$ through $\Psi(\xi)\simeq\xi^{1/\delta}$ for $\xi\gg 1$ (and thus $\delta$ in Eq. (7) corresponds to the critical exponent $\delta$). The scaling exponents $\beta$ and $\Delta$ in Eq. (2) correspond to $\beta$ and $\Delta$ in Eq. (19), where the exponents $\beta$, $\Delta$, and $\delta$ in critical phenomena satisfy Eq. (10) as shown in Appendix B.
In other words, the critical exponents $\beta$, $\Delta$, and $\delta$ in the present case are the ones summarized in Tab. 1. This table suggests that these three exponents are independent, where Eq. (10) is more violated as $e$ increases, and that that the exponents are monotonic functions of the thickness $e$. In addition, we may expect from the above arguments on the universality of $\Gamma(\xi)$, which defines one of the exponents $\delta$, the exponents are universally shared by the systems with different $R$ (and $\eta$) for a given $e$ (at $D_{0}=1$ mm). Namely, the present universality class may be determined by a continuous parameter $e$. This is in contrast with typical critical phenomena, in which universality class is determined by discrete parameters such as $n$ and $d$, although examples of critical exponents depending on continuous parameters are known when marginal operators play a role, which includes non-linear diffusion equations [22].
@@ IV Conclusion
We investigate the pre-detachment dynamics of a bubble surrounded by a viscous liquid in a confined geometry focusing on the dependence on the lubricating film thickness $e$ for a fixed $D_{0}$ (=1.0=1.0 mm). The present results are in contrast with our previous studies on the pre-detachment dynamics [16, 17] in that (I) we add a third regime, the corn-forming regime with breakup, to the previously found two regimes, the sheet-forming regime with breakup and the corn-forming regime without breakup, and that (II) we find the scaling exponents characterizing the self-similarity are dependent on a length. We elucidate similarity between the present singular dynamics and critical phenomena, showing that the counterpart of the critical exponents are dependent on a continuous variable $e$ and could be shared by a universality class formed by the systems with different $R$ (and $\eta$) for a fixed $e$ (at $D_{0}=1.0$ mm). The dependence on the continuous variable suggest the existence of an uncountably infinite number of universality classes. The physical origin of the dependence on $e$ but not on $R$ could be understood from separation of scales, as explored in the study of the post-breakup [18].
The analogy between the singular dynamics and critical phenomena elucidated in the present study underscores the importance of exploring symmetry in the future study of the singular dynamics by exploiting confined geometries for the following reasons. (1) In critical phenomena, which has propelled the development of the modern physics from soft and hard condensed matter to non-equilibrium systems and active matter [23, 24, 25, 26], dimensionality and symmetry have played a vital role in its development. (2) The singular dynamics is widely observed in nature and a variety of confined geometries are increasingly important in industrial applications and natural phenomena, from microfluidics [27, 28], petroleum industry [29], to geology [30]. In fact, it is recently shown that the analogy with critical phenomena can be deepened further by developing a renormalization group analysis for the breakup of a bubble in the non-confined geometry [31], which will be further explored elsewhere for the present case.
Figure 6: Snapshots obtained at 2000 fps near the detachment point for $e=0.5$ mm (top) and $e=1.0$ mm (bottom) at $t-t_{d}=-1.5$ to +1.5+1.5 ms, respectively. The snapshots at the center correspond to $t-t_{d}=-0$ ms, where ”the order parameter” $h_{m}$ is not precisely zero.
@@ Acknowledgements.
Appendix A Dynamics close to the breakup time
Appendix B Scaling relations in critical phenomena